OpenStudy (anonymous):

which property justifies the following equation:7(0+5)=7.0+7.5

Parth (parthkohli):

It starts with D, It ends with E.

OpenStudy (anonymous):

dist.

OpenStudy (anonymous):

distributive property of multiplication over addition

Parth (parthkohli):

\[\rm a(b+c) = a\cdot b + a\cdot c\]